# SafeVLA: Towards Safety Alignment of Vision-Language-Action Model via Safe Reinforcement Learning

### v1

@@ 3 Problem Formulation
$\mathcal{J}^{c_{i}}(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}\left(s_{t},s_{t+1},a_{t}\right)\right],$ (2)
denoting the expected cumulative cost associated with constraint $i$. The set of feasible policies is then defined as
$\Pi_{\mathcal{C}}=\bigcap_{i=1}^{m}\left\{\pi_{\theta}\in\Pi_{\Theta}\mid\mathcal{J}^{c_{i}}(\pi_{\theta})\leq b_{i}\right\},$ (3)
which represents the set of policies that satisfy all safety constraints. Our objective is to determine an optimal policy that adheres to these constraints while following any instruction $l\in\mathcal{L}$. Formally, we aim to solve
$\pi^{*}=\arg\max_{\pi_{\theta}\in\Pi_{\mathcal{C}}}\mathcal{J}(\pi_{\theta}).$ (4)
4 Method: SafeVLA
As shown in Figure 1, we performed simulation experiments in AI2THOR (Kolve et al., 2017) using a highly diverse and realistic set of houses (Deitke et al., 2022) and 3D object assets (Deitke et al., 2023). We identify two types of unsafe behavior exhibited by the vision-language-action (VLA) model. First, the model ignores the safety of objects that are not related to the task, leading to behaviors such as moving, damaging, or misusing hazardous objects. Second, it disregards the safety of the robot’s hardware, resulting in reckless movements and repeated collisions with fixed environmental structures.
To address these challenges, we model these unsafe behaviors within the CMDP framework. Specifically, we specify two constraints: the Object Safety Constraint and the Robot Safety Constraint. Next, we fine-tune SPOC within the SafeVLA framework to integrate these safety constraints into the model.
4.1 Safety Constraints
Our scenes and tasks are derived from the AI2THOR series of works (Kolve et al., 2017), which provide highly realistic 3D indoor environments. By combining scenes generated by ProcTHOR (Deitke et al., 2022) with 3D assets annotated by Objaverse (Deitke et al., 2023), we evaluate the safety of existing models in large-scale and highly diverse scenes. The CHORES tasks (Ehsani et al., 2024), including ObjNav, PickUp, and Fetch, require the model to integrate various capabilities such as natural language reasoning and image understanding, navigation, manipulation, object recognition, and long-term memory. These attributes make the CHORES tasks well-suited for the large-scale evaluation and training of VLAs.
In this context, to assess the effectiveness of SafeVLA, we incorporate our Object Safety Constraint and Robot Safety Constraint into the CHORES tasks, creating the Safety-CHORES tasks: Safety-ObjNav, Safety-PickUp, and Safety-Fetch. We define these constraints using separate cost functions as follows:
$c_{\text{object}}=\begin{cases}2,&\text{if }\exists\,i,j\in I\,(i\neq j)\text{ such that }\\ &\quad\left(\|\Delta d_{i}\|_{\infty}>D\vee\|\Delta\theta_{i}\|_{\infty}>\Theta\right)\\ &\quad\wedge\left(\|\Delta d_{j}\|_{\infty}>D\vee\|\Delta\theta_{j}\|_{\infty}>\Theta\right),\\ 1,&\text{if }\exists\,i\in I\text{ such that }\\ &\left(\|\Delta d_{i}\|_{\infty}>D\vee\|\Delta\theta_{i}\|_{\infty}>\Theta\right)\\ &\quad\wedge\forall\,j\in I\setminus\{i\},\\ &\left(\|\Delta d_{j}\|_{\infty}\leq D\wedge\|\Delta\theta_{j}\|_{\infty}\leq\Theta\right),\\ 0,&\text{otherwise},\end{cases}$
where $I$ is the set of indices corresponding to the objects under consideration, and $\Delta d_{i}\in\mathbb{R}^{3}$ and $\Delta\theta_{i}\in\mathbb{R}^{3}$ represent the changes in position and orientation, respectively, of the $i$-th object between time steps $t$ and $t+1$. The constants $D$ and $\Theta$ denote the threshold values for position and orientation changes, which trigger the corresponding cost.
$c_{\text{robot}}=\begin{cases}1,&\text{if }S_{\text{collided}}\cap S_{\text{forbidden}}\neq\emptyset,\\ 0,&\text{otherwise,}\end{cases}$
where $S_{\text{forbidden}}$ is the set of objects in the environment whose collision would damage the robot’s hardware, and $S_{\text{collided}}$ is the set of objects that collide with the robot at the current time step $t$.
4.2 Safe Reinforcement Learning
We formulate the optimization objective of the VLA model $\pi_{\theta}(\cdot\mid l,\mathcal{H}_{t})$ within the constrained Markov decision process (CMDP) framework as follows:
$\begin{split}\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta},l\sim L}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}(l,s_{t})\right],\\ \text{s.t.}\quad\forall i,\sum_{t=0}^{\infty}\gamma^{t}c_{i}(s_{t},s_{t+1},a_{t})\leq 0,\end{split}$ (5)
where $\gamma\in[0,1)$ is the discount factor, $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory and $\tau\sim\pi_{\theta}$ denotes the trajectory distribution dependent on $\pi_{\theta}$: $s_{0}\sim\mu_{0}$, $a_{t}\sim\pi_{\theta}(\cdot|l,\mathcal{H}_{t})$, $s_{t+1}\sim\mathbb{P}(\cdot|s_{t},a_{t})$. At each time step $t$, the policy considers a temporal context window defined by $\mathcal{H}_{t}=\{(s_{t-n},a_{t-n}),(s_{t-n+1},a_{t-n+1}),\dots,(s_{t-1},a_{t-1}),s_{t}\}$, which contains the history of the past $n$ state-action pairs along with the current state $s_{t}$. This equation expresses our primary objective: to maximize the expected reward derived from the VLA model’s behavior while ensuring that all generated trajectories comply with the safety constraints.
However, directly solving equation (5) through reinforcement learning (RL) is not feasible. Therefore, we recast the safety constraints in an expectation form, aligning them with the reward optimization objective, and introduce hyperparameters $b_{i}$ to regulate the probability that the constraints $c_{i}$ are satisfied. The objective function $\mathcal{J}_{r}(\theta)$ and constraint function $\mathcal{J}_{c_{i}}(\theta)$ are defined as follows:
$\mathcal{J}_{r}(\theta)$ $\triangleq\mathbb{E}_{\tau\sim\pi_{\theta},l\sim L}\left[\sum_{t=0}^{\infty}r(l,s_{t})\right],$ (6)
$\mathcal{J}_{c_{i}}(\theta)$ $\triangleq\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}c_{i}(s_{t},s_{t+1},a_{t})\right]+b_{i}.$ (7)
Then, the surrogate objective corresponding to the original problem is given by
$\begin{split}&\max_{\theta}\mathcal{J}_{r}(\theta),\\ ~{}~{}~{}\text{s.t.}&~{}~{}\forall i,\mathcal{J}_{c_{i}}(\theta)\leq 0.\end{split}$ (8)
To address this constrained optimization problem, we apply the Lagrangian method, which enables us to find the extremum of the objective function over the feasible set defined by the constraints. Specifically, we reformulate the primal problem in equation (8) into its unconstrained Lagrangian dual form:
$\min_{\theta}\max_{\lambda\geq 0}\left[-\mathcal{J}_{r}(\theta)+\sum_{i=0}^{n}\lambda_{i}\mathcal{J}_{c_{i}}(\theta)\right],$ (9)
where $\lambda\geq 0$ are the Lagrange multipliers and $n$ is the number of constraints. The Lagrangian dual form is mathematically equivalent to the primal problem, with the Lagrange multipliers representing the relationship between the gradients of the constraint functions and the objective function in both magnitude and direction. In the optimization process, we alternate between updating the VLA model parameters $\theta$ and the Lagrange multipliers $\lambda$, iteratively solving the min-max problem in equation (9) (for further details, refer to Appendix B.1). This approach ensures that the magnitude of the Lagrange multiplier dynamically reflects the safety risk of the updated model, thus maintaining a balance between safety and task performance throughout the optimization process and preventing excessive optimization of one objective to the detriment of the other.
Table 1: Performance comparison across methods. This table presents a comprehensive comparison of SafeVLA with other methods across three tasks: Safety-ObjNav, Safety-PickUp, and Safety-Fetch. For the penalty in reward shaping, we set it to 1 and select the 20% converged cumulative cost corresponding to the task in FLaRe as the threshold $b_{i}$ for $c_{i}$. The performance metrics include success rate (SR), cumulative robot cost (RC), and cumulative object cost (OC), where higher SR values and lower RC and OC values are desirable. SafeVLA outperforms existing methods in both safety and task performance. It delivers superior safety outcomes and more effectively balances the two optimization objectives compared to other imitation learning (IL) + reinforcement learning (RL) approaches. Moreover, relative to the IL and RL-only baselines, SafeVLA demonstrates overwhelming advantages in both safety and task performance. The orange background of the rows indicates the type of method using privileged information and the bold text indicates the best method per column.
@@ 4.2 Safe Reinforcement Learning
Type Methods SR $\uparrow$ RC $\downarrow$ OC $\downarrow$ SR $\uparrow$ RC $\downarrow$ OC $\downarrow$ SR $\uparrow$ RC $\downarrow$ OC $\downarrow$
IL+RL SafeVLA 0.865 1.698 0.156 0.928 0.246 0.126 0.637 7.93 0.154
FLaRe 0.822 12.021 0.335 0.912 5.526 1.55 0.605 41.407 1.957
FLaRe-RS 0.75 4.42 0.335 0.918 7.216 0.28 0.45 16.16 2.03
IL SPOC-DINOv2 0.43 13.199 0.305 0.86 7.678 2.61 0.14 10.68 3.29
SPOC-SigLip-S 0.584 14.427 0.191 0.883 2.193 3.918 0.14 29.116 3.297
SPOC-SigLip-L 0.38 17.32 0.274 0.83 2.029 3.684 0.135 38.169 3.222
SPOC-SigLip-S w/GT det 0.815 22.715 0.829 0.9 8.641 5.271 0.597 35.064 5.05
SPOC-SigLip-L w/GT det 0.849 16.904 0.593 0.918 1.064 2.824 0.561 24.105 2.502
RL-Only Poliformer 0.804 6.6 2.618 N/A N/A N/A N/A N/A N/A
@@ 5 Experiments
In this section, we demonstrate the effectiveness of SafeVLA in enhancing both model safety and task performance. In particular, we address the following questions:
  • What benefits arise from decoupling safety and task performance objectives in SafeVLA? (Section 5.2.1)
  • Does SafeVLA balance safety and task performance efficiently? (Section 5.2.2)
  • Can SafeVLA generalize to out-of-distribution (OOD) perturbations? (Section 5.2.2)
5.1 Experiment Setup
Evaluation, Tasks and Models
The AI2THOR simulator (Kolve et al., 2017), combined with 150K houses generated by the ProcTHOR pipeline (Deitke et al., 2022) and over 800K 3D object assets annotated by Objaverse (Deitke et al., 2023), provides highly realistic environments in terms of rendering quality and scene complexity. Leveraging these rich and diverse settings, we validate the effectiveness of SafeVLA across a spectrum of challenging tasks.
In the Safety-ObjNav tasks, the robot must navigate through multiple rooms to find a designated object. In contrast, the Safety-PickUp tasks start with the robot in a fixed position, instructing it to pick up a specific object using its arm and gripper. The Safety-Fetch tasks combine both challenges, requiring the robot to first navigate across multiple rooms to locate the target object and then pick it up. For all tasks, natural language instructions are provided, and the robot’s observations consist solely of RGB images captured from two egocentric cameras positioned at different angles to facilitate both navigation and manipulation.
Beyond the defined safety constraints, these tasks present additional challenges such as image-semantic fusion, object recognition, long-term planning, navigation, manipulation, and coping with highly diverse scenes and objects. These complexities thoroughly test the vision-language-action (VLA) model’s safety and capabilities. In response to these challenges, we adopt SPOC as the base VLA model for fine-tuning. SPOC is trained on millions of frames of expert trajectories and has demonstrated impressive performance in both simulation and real-world deployments. For detailed information about task definitions, model observations, and safety constraints, please refer to Appendix C.
Evaluation Metrics
We evaluate the SafeVLA-trained model on separate test sets for each of the three tasks. Borrowing from traditional safety considerations in robotics (Lozano-Pérez & Kaelbling, 2014; Castillo-Lopez et al., 2020), our evaluation focuses on two primary metrics: the task success rate and the cumulative cost incurred in environments. As detailed in Section 4.1, the cumulative cost metric comprises two distinct components: (1) cumulative object cost, which quantifies the risk of impacts on external objects, and (2) cumulative robot cost, which assesses potential damage to the robotic hardware. This dual-cost formulation facilitates a comprehensive safety evaluation by simultaneously addressing both environmental preservation and robotic self-protection objectives. All experiments are conducted using the same hyperparameter configurations as specified in Appendix B.2.
Figure 2: Cumulative cost distribution analysis. Left: Distribution of cumulative cost across robot trajectories in the test set after fine-tuning with SafeVLA and FLaRe. The cost distribution shows a marked improvement in SafeVLA, eliminating high-risk unsafe behaviors (cumulative cost >10). Middle: Cumulative cost distribution when the task succeeds highlights the minimal severity and frequency of unsafe behaviors in SafeVLA compared to FLaRe. Right: Cumulative cost distribution when the task fails, where SafeVLA maintains a consistent cost distribution regardless of success, emphasizing its decoupling of safety and task performance objectives.
Table 2: OOD results across tasks. This table presents the results of evaluating SafeVLA across various OOD perturbations. The table compares SafeVLA’s performance in three task categories: Safety-ObjNav, Safety-PickUp, and Safety-Fetch, under different types of disturbances, including color, light, material, and all combined. Performance metrics include success rate (SR), cumulative robot cost (RC), and cumulative object cost (OC), where higher SR values and lower RC and OC values are preferred. The results demonstrate that SafeVLA maintains superior safety and performance in the presence of OOD perturbations. The average performance changes for each task are also shown, highlighting the robustness of SafeVLA in adapting to various challenges. The bold text indicates the best method per column.

### v2

@@ 3 Problem Formulation
$\Pi_{\mathcal{C}}=\left\{\pi_{\bm{\theta}}\in\Pi_{\Theta}\mid\mathbb{E}_{\pi_{\bm{\theta}}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}(s_{t},a_{t})\right]\leq b_{i},\forall i=1,\dots,m\right\}.$ (1)
Formally, we aim to solve
$\pi^{*}=\arg\max_{\pi_{\bm{\theta}}\in\Pi_{\mathcal{C}}}\mathcal{J}(\pi_{\bm{\theta}}).$ (2)
4 Implementing the Integrated Safety Approach
We argue that VLA safety requires an integrated safety approach (ISA), rather than a single method. Specifically, an ISA address four interconnected aspects, as shown in Figure 1: (i) modeling safety-critical aspects of tasks and environments; (ii) eliciting latent and diverse unsafe behaviors from existing policies; (iii) constraining the VLA’s learning process to integrate these safety considerations; and (iv) assuring the resulting model’s safety through rigorous and targeted evaluation. In this section, we present our methodologies into each of these aspects.
4.1 Modeling Safety: Scenes, Specifications, and Tasks
In our investigation, we focus on a mobile manipulation setting. The static part of each task $t_{i}\in T$ is defined as $(e_{i},x_{i},q_{i},G,\Phi,\Psi)$. Here, $e_{i}\in E$ is the scene, $x_{i}$ and $q_{i}$ are the randomly selected initial robot position and orientation, $G$ is the set of object categories in $e_{i}$, $\Phi$ is a set of state-action safety predicates, and $\Psi$ is a set of trajectory-level safety predicates. A safety predicate serves as a compact representation for identifying unsafe behaviors. It can be expressed as either a state-action predicate $\phi:\mathcal{S}\times\mathcal{A}\to\{0,1\}$ or a trajectory-level predicate $\psi:\mathcal{H}\to\{0,1\}$.
Each state-action predicate is defined using compositional logic:
$\phi(s,a)=1\iff P_{s}(s)\land P_{a}(a)\land R(s,a),$
where $P_{s}$ and $P_{a}$ capture relevant conditions on states and actions, and $R$ represents the risk-inducing relation. Similarly, trajectory predicates are defined as:
$\psi(\tau)=1\iff\exists t_{0},\dots,t_{k}\in[0,\text{len}(\tau)]\text{ s.t. }\left(\bigwedge_{i=0}^{k}E_{i}(s_{t_{i}},a_{t_{i}})\right)\land R_{\text{temporal}}(\{(t_{j},s_{t_{j}},a_{t_{j}})\}_{j=0}^{k},\tau),$
where each $E_{i}(s_{t_{i}},a_{t_{i}})$ is an event predicate that evaluates to true if a specific condition holds for the state-action pair $(s_{t_{i}},a_{t_{i}})$ at time $t_{i}$, and $R_{\text{temporal}}(\cdot)$ is a predicate describing the temporal structure.
To instantiate $t_{i}$, we dynamically augment the static components with a natural language instruction $l$. Specifically, we randomly select an object category $g\in G$ as the goal, and then sample a natural language instruction $l$ to specify $g$. Inspired by ehsani2024spoc , we build three categories of tasks:
  • Safety-ObjNav: The robot must navigate through multiple rooms to locate a designated object.
  • Safety-PickUp: The robot begins in front of a surface and is instructed to pick up a specific object.
  • Safety-Fetch: This task requires the robot first navigate to find the target object and then pick it up.
4.2 Eliciting Risks: Uncovering Latent Unsafe Behaviors
To ensure comprehensive risk elicitation and to prevent policies from overfitting to limited scenarios, maximizing the diversity of both environmental settings and interactable objects is critical. Therefore, we utilize a large-scale dataset of 150K diverse indoor scenes generated by ProcTHOR deitke2022️ , alongside Objaverse deitke2023objaverse , which provides an extensive library of 800K 3D assets. The simulation is conducted in the AI2THOR kolve2017ai2 simulator, which supports photo-realistic rendering quality, object state changes, arm-based manipulation, and causal interactions.
Building upon this foundation of diverse scenes and objects, to further systematize risk elicitation and ensure targeted coverage of known problematic scenarios, we identify and leverage several safety critical components. These are not separate entities but rather specific environmental features (e.g., narrow corners) or challenging object arrangements (e.g., fragile collections) that are instantiated or frequently occur within the aforementioned large-scale scenes. The safety critical components considered in our study include (see Appendix D for details):
Figure 2: Upper: Conceptual diagrams of each safety critical component. Lower: Corresponding photorealistic examples from our simulation environment.
  • Corners ($\phi_{\text{corner}}$): Situations where navigation into confined spaces like narrow corners leads to the robot becoming stuck or incurring repeated collisions.
  • Blind Spots ($\psi_{\text{blind spot}}$): Collisions with previously seen but currently unobserved obstacles due to failures in maintaining short-term spatial awareness.
  • Fragile Collections ($\psi_{\text{fragile collection}}$): Scenarios involving collateral damage to nearby fragile items during manipulation tasks, often due to object density or precarious placements.
  • Critical Points ($\psi_{\text{critical point}}$): Incidents where robot actions, even indirect ones, destabilize precariously positioned objects (e.g., a knife on an edge), causing them to fall.
  • Dangerous Equipment ($\phi_{\text{dangerous equipment}}$): Prohibited interactions with intrinsically hazardous objects like active stovetops or exposed wiring, which demand strict avoidance.
By incorporating these diverse scenes, objects, and safety critical components, we propose Safety-CHORES to systematically elicit a wide spectrum of potential safety violations, thereby generating rich, safety-aware data. These complex tasks require VLAs to integrate natural language understanding, visual reasoning, and long-horizon planning, while adhering to the modeled safety constraints.
4.3 Constraining Policies: Safe Reinforcement Learning for Alignment
Once safety specifications are modeled and data of potential risks can be elicited, we leverage SafeRL techniques to effectively integrate these safety considerations into the VLA’s policy learning process.
A preliminary step is translating safety predicates ($\phi,\psi$) into cost signals for the cost-returns $\mathcal{J}_{c_{i}}(\bm{\theta})$. State-action predicate ($\phi_{k}$) violations incur a cost of 1 at the violating timestep $t$, otherwise 0. For trajectory-level predicates ($\psi_{j}$), a cost of 1 is attributed solely to the final step of the violating segment in this initial exploration. The credit assignment for $\psi_{j}$ remains an area for exploration in future work. The Lagrangian method is a general solution for SafeRL. By employing the Lagrangian relaxation technique (NoceWrig06, ), Equation 2 is transformed into an unconstrained safe optimization problem:
$\min_{\bm{\theta}}\max_{\lambda\geq 0}[-\mathcal{J}_{r}(\bm{\theta})+\sum_{i=0}^{n}\lambda_{i}\mathcal{J}_{c_{i}}(\bm{\theta})],$ (3)
where $\lambda_{i}\geq 0$ is the Lagrange multiplier and $n$ is the number of constraints.
Solving the min-max optimization in Equation 3 necessitates an iterative refinement process, where updates to the VLA model parameters $\bm{\theta}$ are interleaved with those to the Lagrange multipliers $\lambda$. It optimizes for safety first, then maximizing task performance. This trade-off ensures that the VLA model adheres to safety requirements while maximizing task performance within these constraints.
4.4 Safety Assurance: Evaluating Aligned VLAs
The final aspect of ISA is the assurance of safety through comprehensive evaluation. Our assurance methodology systematically assesses the model’s safety performance across several dimensions:
  • Test-time Safety evaluates the model’s adherence of safety constraints through performance on held-out test sets and out-of-distribution (OOD) perturbations. The primary goal is to quantify the learned safe behaviors in the training phase.
  • Long-tail Safety considers the model’s safety on statistically infrequent events. Ensuring that the model does not exhibit long-tail safety issues is crucial for robust safety in real-world deployments.
  • Extreme Failure Safety focuses on the model’s safety and behavior to catastrophic failures. This is particularly assessed in situations where task completion may be impossible.
Table 1: Performance comparison across methods. The orange background of the rows indicates the methods using privileged information and the bold text indicates the best method per column.
@@ 4.4 Safety Assurance: Evaluating Aligned VLAs
Type Methods SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$
IL+RL ISA 0.865 1.854 0.928 0.372 0.637 8.084
FLaRe 0.822 12.356 0.912 7.076 0.605 43.364
FLaRe-RS 0.75 4.755 0.918 7.496 0.45 18.19
IL SPOC-DINOv2 0.43 13.504 0.86 10.288 0.14 13.97
SPOC-SigLip-S 0.584 14.618 0.883 6.111 0.14 32.413
SPOC-SigLip-L 0.38 17.594 0.83 5.713 0.135 41.391
SPOC-SigLip-S w/GT det 0.815 23.544 0.9 13.912 0.597 40.114
SPOC-SigLip-L w/GT det 0.849 17.497 0.918 3.888 0.561 26.607
RL-Only Poliformer 0.804 9.218 N/A N/A N/A N/A
Figure 3: Cumulative cost distribution analysis. Left: Distribution of cumulative cost across robot trajectories in the test set after fine-tuning with ISA and FLaRe. Middle: Cumulative cost distribution when the task succeeds. Right: Cumulative cost distribution when the task fails.
@@ 5 Experiments
In this section, we aim to answer the following questions: (I) Can ISA outperform standard VLA fine-tuning methods? (§ 5.2.1); (II) How do ISA-aligned VLAs qualitatively handle risks and failures? (§ 5.2.2); (III) Which components within ISA critically impact its safety-performance balance? (§ 5.2.3) (IV) Do learned safety behaviors generalize to OOD scenarios and extreme failures? (§ 5.2.4)
5.1 Experimental Setup
Tasks and Environments.  Our primary experiments utilize Safety-CHORES. To contextualize the unique challenges posed by Safety-CHORES, we also conduct comparisons on other benchmarks kolve2017ai2 ; deitke2020robothor ; deitke2022️ focusing on object navigation and generally lack the safety features of Safety-CHORES.
Baseline Methods.  We compare ISA against a comprehensive set of baselines that represent various paradigms for VLA training and fine-tuning. IL-only: SPOC ehsani2024spoc , which is a state-of-the-art imitation learning method. IL-only (Ground Truth): SPOC augmented with ground truth information. These models can thoroughly showcase the potential upper bound of IL methods. IL+RL (Standard): FLaRe hu2024flare , which fine-tunes pre-trained VLAs using reinforcement learning focused solely on task performance. IL+RL (Reward Shaping): FLaRe-RS, a variant of FLaRe where safety costs are directly used as penalties on reward, representing a common heuristic for addressing safety. RL-Only: Poliformer zeng2024poliformer , an end-to-end RL approach for navigation tasks.
Initial IL Model.  We begin our experiments with the SPOC-DINOv2 model. We select it as our initial model for two main reasons. First, SPOC is a state-of-the-art VLA trained solely on simulated data. Second, it demonstrates strong transferability to real-world deployment, making it suitable for safety-critical data collection. We also evaluate ISA on other VLA models (i.e., EmbCLIP khandelwal2022simple , Embodied-Codebook eftekhar2023selective and their variants with different vision encoders).
Evaluation Metrics.  Borrowing from safety considerations in robotics lozano2014constraint ; castillo2020real , our evaluation focuses on two metrics: the task success rate (SR) and the cumulative cost (CC). The CC is an aggregate measure of all safety violations throughout an episode. For a trajectory $\tau$ of length $L$ and $K$ distinct safety constraint types, it is computed as $CC(\tau)=\sum_{k=1}^{K}\sum_{t=0}^{L-1}c_{k}(s_{t},a_{t})$, where $c_{k}(\cdot)$ is the cost incurred from violating the $k$-th safety constraint at step $t$.
Figure 4: Effectiveness of ISA across diverse VLA models and benchmarks.
5.2 Main Results
5.2.1 Comparative Performance: ISA vs. Standard Methods
We first evaluate the effectiveness of ISA in enhancing VLA safety while preserving task performance. In Table 1, we present the performance of ISA against baseline methods on Safety-CHORES. ISA demonstrates substantial safety improvements, achieving an average reduction in CC of 83.58% compared to the strongest task-focused RL baseline, FLaRe. This significant decrease is consistent across all tasks, as illustrated by per-room safety improvements in Figure 9. Crucially, these safety enhancements are accompanied by maintained task performance. ISA achieves an average SR increase of 3.85% compared to FLaRe, outperforming IL-only baselines and matching or exceeding other RL-based methods. This indicates ISA effectively trades off the safety and task performance, in contrast to approaches that solely optimize for task performance.
5.2.2 Qualitative Insights: Risk Handling and Failure Modes
In Figure 3 (Left), we present the distribution of cumulative safety costs for ISA and FLaRe across all test trajectories. A key observation is that ISA eliminates trajectories with extremely high safety costs (cumulative cost >10). The upper bound of unsafe behavior severity in ISA is reduced to 1/35th of that in FLaRe, indicating a significant mitigation of catastrophic safety failures. This shift in distribution demonstrates ISA’s effectiveness in mitigating long-tail risks, where a small number of trajectories could otherwise account for a disproportionate amount of unsafe behaviors.
Further analysis, shown in Figure 3 (Middle and Right), reveals a difference in how safety correlates with task success. For FLaRe, higher safety costs are more prevalent in task failures, suggesting that unsafe behaviors often contribute to or coincide with failure. Logistic regression and Pearson correlation tests (see Appendix A for more details) confirm a significant negative correlation between cost and success for FLaRe (p < 0.01). In contrast, ISA exhibits a more consistent cost distribution regardless of task outcome. The T-test rejects the correlation for ISA, indicating that the learned safety paradigm is largely decoupled from task success. Even when ISA fails a task, it tends to do so more safely, avoiding safety violations. This suggests a deeper integration of safety principles rather than superficial avoidance. For further cases and behavior analysis, please refer to Appendix B.1.
Figure 5: Comparative performance of VLA models on multiple benchmarks. Left: SR of each model per benchmark. Right: CC incurred by each model on these benchmarks.
5.2.3 Ablation Studies: Impact of Key ISA Design Choices
To understand the contribution of specific design choices in ISA, we conduct several ablation studies.
Figure 6: ISA with fixed penalty coefficients.
Importance of Risk Elicitation.  The importance of risk elicitation is demonstrated by an ablation study in Figure 7 (Left). When the standard ISA training recipe was applied to simplified one-room scenes without safety critical components, safety performance degraded considerably. This ablated model yielded a CC nearly three times higher than the full ISA’s (5.01 vs. 1.854) and even performed worse than the FLaRe-RS baseline, alongside a reduced SR (0.645 vs. 0.865). This significant decline, particularly in safety despite identical constraining mechanisms, underscores that rich elicitation environments are indispensable for achieving safety alignment superior to heuristic approaches.
ISA Generalizability to Different VLA Models.  In Figure 4, we validate the generalizability by applying ISA’s alignment process to several distinct VLA base models. The results consistently show that ISA alignment leads to substantial improvements across these models, evidenced by significant reductions in CC alongside stable SR when evaluated on Safety-CHORES and other benchmarks.
Safety Challenges Posed by Safety-CHORES.  In Figure 5, we demonstrate the applicability of Safety-CHORES to various VLA models and observe a consistent trend: across various VLA models, the CC on Safety-CHORES (green segments) often more than 2 times that on benchmarks like iTHOR or ProcTHOR. This pronounced difference is observed under identical safety evaluation mechanisms applied to all benchmarks; however, standard benchmarks inherently lack the safety-critical environmental designs.
Figure 7: Left: Ablation of the risk elicitation component. Middle: Ablation on cost thresholds $b_{i}$. Right: Safety in extreme failure scenarios.
Importance of Lagrangian Multipliers.  The Lagrangian dual formulation (Equation 3) uses dynamic multipliers $\lambda$ to balance reward and cost objectives. We compare this against baselines using fixed penalty coefficients for safety costs, as shown in Figure 6. The results demonstrate that our approach with dynamic Lagrangian multipliers achieves a superior trade-off, adhering to the cost limit while attaining a higher success rate than any fixed-penalty baseline that meets the same cost constraint. This highlights the benefit of the adaptive constraining mechanism provided by the Lagrangian method for effectively balancing safety and task performance.
Impact of Cost Threshold $b_{i}$ The choice of the safety cost threshold $b_{i}$ in the CMDP formulation (Equation 1) directly influences the strictness of the safety constraints. In Figure 7 (Middle), we shows the performance on Safety-ObjNav when varying $b_{i}$ (e.g., 10%, 20%, 50% of FLaRe’s converged cumulative cost 11.5982). As observed, stricter thresholds lead to lower realized safety costs, demonstrating effective constraint enforcement. However, excessively strict thresholds (e.g., 10%) might slightly impact SR. The chosen 20% threshold offers a balance.
5.2.4 Robustness: Generalization to OOD Scenarios and Extreme Failures
OOD Perturbation Results.  In Table 2, we presents the performance of ISA on Safety-CHORES tasks under four types of OOD perturbations: color, lighting, material, and all combined. The average changes reported at the bottom of Table 2 indicate that, the safety benefits and reasonable task performance achieved by ISA are largely preserved under OOD challenge. For example, in Safety-ObjNav, while SR sees a modest average decrease of 0.042 under OOD conditions, the safety metrics remain significantly better than those of unaligned baselines in standard conditions. The impact of perturbations on safety is limited across all tasks; safety costs generally remain contained and highly stable, with instances like PickUp +All even showing a decrease in CC. This robustness indicates that the learned safe behaviors are not superficial. (see Appendix E.4 for OOD setup.)
Table 2: OOD results across tasks.